In a poisson frequency distribution, frequency corresponding to 3 successes in 2/3 times frequency corresponding to 4 successes. Find the mean and standard deviation. 


1
Expert's answer
2021-03-16T14:20:33-0400

Denote a number of successes as "k" .

Then probability for Poisson distribution is

"P(k)=\dfrac{\lambda^k\cdot e^\lambda}{k!}"

Mean and variance are both equal to "\lambda" .

So,

mean "\bold{E}[X]=\lambda"

standard deviation "\sigma=\sqrt{\bold{Var}[X]}=\sqrt{\lambda}" .


"P(3)=\dfrac{\lambda^3\cdot e^\lambda}{3!}"


"P(4)=\dfrac{\lambda^4\cdot e^\lambda}{4!}"


"P(3)=\dfrac{2}{3}P(4)"


"\dfrac{\lambda^3\cdot e^\lambda}{3!}=\dfrac{2}{3}\cdot\dfrac{\lambda^4\cdot e^\lambda}{4!}"


"\dfrac{\lambda^3}{3!}=\dfrac{2}{3}\cdot\dfrac{\lambda^4}{4!}"


"\dfrac{1}{3!}=\dfrac{2}{3}\cdot\dfrac{\lambda}{4!}"


"\dfrac{\lambda}{4!}=\dfrac{3}{2}\cdot\dfrac{1}{3!}"


"\lambda=\dfrac{3}{2}\cdot\dfrac{4!}{3!}"


"\lambda=\dfrac{3}{2}\cdot\dfrac{3!\cdot4}{3!}=3\cdot2=6"


Answer

Mean "\bold{E}[X]=6"

Standard deviation "\sigma=\sqrt{6}"